OpenStudy (anonymous):

sovle the following inequality 10x^2-6+7x

OpenStudy (anonymous):

\[10x ^{2}+7x-6 = 10x ^{2} -5x + 12x -6 = 5x(2x-1) + 6(2x-1) = (2x-1)(5x+6)\]

OpenStudy (chihiroasleaf):

there is no inequality sign in your problem :) are you asking to factorize the expression?

OpenStudy (anonymous):

yes factor